Please be as detailed as you can in your steps :) thank you!The next three questions pertain to the situation described below An ice cube slides without friction down a track as shown i
Answer #1

(5) The speed of the cube at the point B 3 mo? = mg (h – hy) v2 = 28(1, - hy) v= 2(9.8)(4.5 – 2.7) = 5.94 m/s (6) The speed o